### v4

# BIS Reasoning 1.0: The First Large-Scale Japanese Benchmark for Belief-Inconsistent Syllogistic Reasoning

Abstract
We present BIS Reasoning 1.0, the first large-scale Japanese dataset of syllogistic reasoning problems explicitly designed to evaluate belief-inconsistent reasoning in large language models (LLMs). Unlike prior datasets such as NeuBAROCO and JFLD, which focus on general or belief-aligned reasoning, BIS Reasoning 1.0 introduces logically valid yet belief-inconsistent syllogisms to uncover reasoning biases in LLMs trained on human-aligned corpora. We benchmark state-of-the-art models—including GPT models, Claude models, and leading Japanese LLMs—revealing significant variance in performance, with GPT-4o achieving 79.54% accuracy. Our analysis identifies critical weaknesses in current LLMs when handling logically valid but belief-conflicting inputs. These findings have important implications for deploying LLMs in high-stakes domains such as law, healthcare, and scientific literature, where truth must override intuitive belief to ensure integrity and safety.
BIS Reasoning 1.0: The First Large-Scale Japanese Benchmark for Belief-Inconsistent Syllogistic Reasoning
1 Introduction
Large language models (LLMs) have demonstrated remarkable performance on various natural language tasks, yet ensuring reliable logical reasoning remains an open challenge Morishita et al. (2024); Nguyen et al. (2023). This issue is particularly critical in high-stakes domains such as law, healthcare, and scientific research, where even subtle reasoning errors can lead to severe consequences Yan et al. (2025). Alarmingly, recent studies show that the persuasive fluency of models like GPT-4 can deceive users into trusting incorrect conclusions Bajpai et al. (2024). Hence, rigorously evaluating and enhancing LLM reasoning accuracy is crucial before deploying them in domains demanding strict logical rigor.
A major concern in current LLM research is their susceptibility to human-like cognitive biases, notably the belief bias—accepting conclusions aligned with prior beliefs regardless of logical validity. This bias poses significant risks in applications requiring impartial logic. For example, Dasgupta et al. demonstrated that LLMs frequently endorse logically invalid arguments simply because their conclusions appear believable, highlighting a critical vulnerability.
Existing benchmarks for evaluating logical reasoning in LLMs exhibit several limitations. Most influential reasoning benchmarks are predominantly English-based Li et al. (2023); Qin et al. (2019); Frohberg and Binder (2022), creating a significant evaluation gap for languages such as Japanese. Although some Japanese datasets exist, like JFLD Morishita et al. (2024), which tests formal logic isolated from real-world knowledge, and NeuBAROCO Ozeki et al. (2024), which covers multiple biases but has limited belief-inconsistent content, no dedicated Japanese-language dataset explicitly targets belief-inconsistent reasoning.
Figure 1: A translated example from the BIS Dataset illustrating a belief-inconsistent syllogism: although the conclusion is logically valid, it contradicts common real-world beliefs.
To address this gap, we introduce BIS Reasoning 1.0—the first Japanese dataset designed explicitly for assessing belief-inconsistent syllogistic reasoning in LLMs. BIS (Belief-Inconsistent Syllogisms) aims to evaluate whether models can uphold logical validity when correct conclusions conflict with typical beliefs or factual knowledge. Figure 1 shows a representative example from the BIS Dataset, illustrating how logically valid conclusions can defy intuitive beliefs. Specifically, our contributions include:
  1. 1.
    BIS Reasoning 1.0 Dataset: We present a carefully curated collection of Japanese syllogistic reasoning problems explicitly constructed to challenge LLMs with logically valid conclusions that contradict common beliefs. This dataset enables the first targeted evaluation of belief-inconsistent reasoning capabilities in Japanese LLMs.
  2. 2.
    Comprehensive Evaluation of Leading LLMs: We benchmark state-of-the-art models—including GPT-4o, GPT-4-turbo, Claude, and prominent Japanese LLMs—under standardized conditions, providing the first systematic comparison of their performance on belief-inconsistent reasoning in Japanese.
  3. 3.
    Detailed Analysis of Performance and Bias: Our analysis identifies significant performance gaps, highlighting that even advanced LLMs struggle disproportionately with belief-inconsistent problems. We quantify these biases, investigate specific syllogistic structures prone to errors, and examine how prompts and chain-of-thought approaches affect reasoning accuracy.
  4. 4.
    Implications for Reliability in Real-World Applications: We discuss critical implications for deploying LLMs in safety-critical domains. BIS Reasoning 1.0 reveals vulnerabilities that standard benchmarks typically overlook, providing crucial insights for improving logical consistency and objectivity in real-world scenarios like law, medicine, and scientific research.
Overall, BIS Reasoning 1.0 contributes significantly to understanding and improving logical reasoning in Japanese-language LLMs. By explicitly evaluating belief-inconsistent reasoning, this work advances efforts toward creating reliable, bias-resistant models suitable for deployment in high-stakes environments.
2 Related Work
Evaluating the logical reasoning abilities of large language models (LLMs) has become a key research focus. Benchmarks like ReClor Yu et al. (2025) and LogiQA Liu et al. (2023) use multiple-choice logic problems derived from standardized exams to test inference beyond surface semantics. Despite advances such as chain-of-thought prompting and neuro-symbolic modeling Kojima et al. (2022), LLMs still struggle to match human performance on tasks requiring rigorous logic.
LLMs not only make logical errors but also exhibit human-like cognitive biases. One well-studied bias is belief bias—the tendency to accept conclusions that align with prior beliefs regardless of logical validity Evans et al. (1983). Studies have shown that LLMs are more accurate on belief-consistent reasoning tasks and frequently misjudge belief-inconsistent syllogisms Ando et al. (2023); Ozeki et al. (2024).
Instruction tuning and RLHF can amplify these tendencies. For instance, models like GPT-4 and Claude, while highly fluent, may reinforce belief-aligned reasoning due to human preferences during fine-tuning Bai et al. (2022). This further underscores the need for benchmarks that reveal latent cognitive biases and test reasoning under belief-conflicting conditions.
In the Japanese language, recent efforts have produced datasets for logical reasoning evaluation, yet limitations remain. JFLD Morishita et al. (2024) focuses on formal deductive reasoning using artificially constructed propositions to isolate logic from world knowledge. While large in scale and diverse in structure, its use of semantically unnatural sentences prevents assessment of reasoning in realistic settings. JaNLI Yanaka and Mineshima (2021) and JAMP Sugimoto et al. (2023) explore adversarial and temporal inference respectively, but they do not test belief-inconsistent reasoning and lack syllogistic structure. NeuBAROCO Ozeki et al. (2024) closely aligns with our goals, exploring belief bias in syllogistic reasoning. However, it falls short as a comprehensive benchmark: the Japanese subset contains fewer than 800 examples for the NLI task and under 100 for the multiple-choice format, and it does not exclusively target belief-inconsistent reasoning. In contrast, BIS Reasoning 1.0 offers a focused and large-scale evaluation specifically designed to test logical robustness under belief-conflicting conditions.
These observations point to a significant gap: current Japanese benchmarks either employ unnatural representations, ignore belief-inconsistent logic, or lack scale and coverage of syllogistic forms. While prior work has identified belief bias in LLMs, comprehensive datasets for evaluating such bias in Japanese syllogistic reasoning remain scarce. BIS Reasoning 1.0 directly addresses this gap by providing a focused, large-scale, and naturalistic benchmark specifically designed to test how LLMs handle logically valid conclusions that conflict with intuitive beliefs.
3 Dataset Construction
We present BIS Reasoning 1.0111https://hf.co/datasets/nguyenthanhasia/BIS_Reasoning_v1.0, a Japanese-language dataset consisting of 5,000 carefully constructed syllogistic reasoning problems. The dataset is specifically designed to test the robustness of logical inference in large language models (LLMs) under conditions of belief inconsistency, where logically valid conclusions explicitly contradict widely held commonsense beliefs. BIS Reasoning 1.0 serves as a diagnostic benchmark for probing whether LLMs can prioritize formal logic over prior-knowledge heuristics in natural language reasoning.
The dataset was developed through a formalized specification process aimed at ensuring both logical rigor and linguistic quality. Each example comprises two premises and one conclusion that is strictly entailed by syllogistic rules, such as classic forms like “All A are B; All C are A; Therefore, All C are B.” Crucially, the conclusions are deliberately chosen to conflict with general knowledge, encouraging model errors driven by belief bias. This design isolates the logical reasoning process from superficial semantic plausibility, exposing potential biases embedded in LLM training data.
Figure 2: Combined category analysis of top raw and final semantic categories in the BIS dataset. The figure illustrates the top 5 raw categories (left), top 5 final categories (right), and the heatmap shows the relationship between these raw and final categories, highlighting category intersections and distribution patterns.
To ensure linguistic fluency and naturalness, all annotators involved in dataset construction were either native Japanese speakers or individuals with advanced Japanese proficiency. Initially, the dataset covers 46 distinct semantic categories (raw categories), ranging from concrete areas like animals, food, and weather to abstract domains such as law, logic, and emotion. These detailed raw categories were subsequently consolidated into 10 broader final categories to facilitate interpretability, ensure topic balance, and support higher-level reasoning analysis. Figure 2 summarizes the distribution and relationship between these raw and final categories, reflecting a concentration in cognitively rich areas valuable for probing belief-inconsistent reasoning in LLMs.
All examples underwent a two-phase quality assurance (QA) process, initially involving manual review of 10% of examples for iterative feedback, followed by comprehensive review to ensure structural validity, language clarity, and semantic diversity. Issues identified included syntactic violations, duplicated content, ambiguous premises, and category imbalance.
4 Experiments
4.1 General Settings
To evaluate how well LLMs handle logically valid yet belief-inconsistent inferences, we formulate BIS Reasoning 1.0 as a diagnostic test focusing strictly on logical judgment. Specifically, models must determine if a given conclusion logically follows from two premises, even when the conclusion contradicts intuitive beliefs.
We framed the task as a binary classification using concise, instruction-based prompts in Japanese, asking models to judge logical entailment by answering “{CJK}UTF8ipxmはい” (Yes) or “{CJK}UTF8ipxmいいえ” (No). Each prompt clearly stated two premises and one conclusion, accompanied by a brief system instruction reinforcing the model’s logical reasoning role. Models were not required to provide explanations, thereby isolating pure logical inference capability from linguistic fluency or explanatory quality.
Accuracy is measured as the proportion of examples for which the model outputs the correct judgment–always “{CJK}UTF8ipxmはい,” since all BIS entries are logically valid. This setup ensures that errors stem from reasoning failures, not linguistic ambiguity or semantic bias.
All models in the experiments (see 4.2) were evaluated under identical zero-shot conditions, using Japanese-language prompts with consistent formatting. No fine-tuning or task-specific adaptation was applied. Evaluation was performed on the full dataset to eliminate sampling variance and enable direct comparison of out-of-the-box reasoning robustness.
4.2 Model Configuration
We evaluated prominent LLMs spanning both general-purpose and Japanese-specialized categories. The general-purpose group included OpenAI’s GPT-4o and GPT-4-turbo, as well as Anthropic’s Claude-3-sonnet and Claude-3-opus. These models are designed for multilingual tasks and are optimized for general reasoning performance across domains. In contrast, the Japanese-specialized models—llm-jp-3-13b, llm-jp-3-13b-instruct3 Aizawa et al. (2024), and stockmark/stockmark-13b [^2] https://hf.co/stockmark/stockmark-13b—were trained or fine-tuned specifically on Japanese data, representing dedicated efforts to advance native Japanese LLM capabilities.
Proprietary models were accessed via official APIs, with consistent configurations including temperature set to zero to ensure deterministic outputs. Open-source models were deployed locally using standardized hardware and inference settings. All prompts were in Japanese, and all 5,000 examples in the dataset were evaluated without sampling. This setup guarantees both fairness and reproducibility across model comparisons.
4.3 Overall Model Performance
Table 1 summarizes the performance of each model on the BIS Reasoning 1.0 dataset, along with complementary results obtained on over 300 belief-inconsistent syllogistic reasoning samples from the NeuBAROCO benchmark. GPT-4o demonstrated the highest accuracy on both datasets, achieving 79.54% on BIS Reasoning 1.0 and notably excelling with 94.01% on NeuBAROCO. GPT-4-turbo and the Japanese-specialized model llm-jp-3-13b showed comparable results, each reaching around 59–60% accuracy on BIS Reasoning 1.0, but significantly lower than GPT-4o. Models specifically fine-tuned on Japanese instructions, such as llm-jp-3-13b-instruct3 and stockmark-13b, exhibited lower performance, around 40–48%, highlighting potential shortcomings in their reasoning robustness. Surprisingly, Anthropic’s Claude models underperformed drastically on the BIS Reasoning 1.0 dataset, with Claude-3-sonnet and Claude-3-opus achieving only 20.34% and 7.18%, respectively, despite relatively strong performances of 78.44% and 61.07% on NeuBAROCO.
Model BIS Reasoning 1.0 NeuBAROCO
OpenAI/GPT-4o 79.54 94.01
llm-jp/llm-jp-3-13b 59.86 67.66
OpenAI/GPT-4-turbo 59.48 67.66
llm-jp/llm-jp-3-13b-instruct3 40.90 38.32
stockmark/stockmark-13b 40.34 47.90
Claude-3-sonnet-20240229 20.34 78.44
Claude-3-opus-20240229 7.18 61.07
Table 1: Accuracy of models on the BIS Reasoning 1.0 dataset and NeuBAROCO belief-inconsistent syllogisms.
These results indicate that superior performance in belief-inconsistent reasoning is heavily influenced by a model’s underlying training approach and inherent architectural reasoning capabilities rather than mere specialization in Japanese or model size alone. Notably, GPT-4o’s ability to consistently achieve high accuracy suggests a strong robustness to belief bias, allowing it to distinguish logical validity effectively despite intuitive contradictions.
The varying performances between datasets also highlight the sensitivity of model evaluations to task formulation and dataset characteristics. The relatively high NeuBAROCO scores of Claude models contrast sharply with their underperformance on BIS Reasoning 1.0, emphasizing the importance of employing diverse benchmarks to comprehensively assess model reasoning capabilities, particularly in challenging belief-inconsistent contexts.
4.4 Detailed Error Analysis for GPT-4o
To further investigate GPT-4o’s behavior on belief-inconsistent syllogisms, we conducted additional experiments on 100 cases where GPT-4o initially failed. These error samples were reassessed using several carefully designed prompts, each emphasizing different reasoning approaches, linguistic styles, or explicit instructions about the belief-inconsistent nature of the task. Figure 3 summarizes the accuracy results across these varied prompts.
Figure 3: Error Sample Accuracy by Prompt Type for GPT-4o
Prompt Impact Analysis
The prompt emphasizing an explicit chain-of-thought (CoT) reasoning strategy yielded the highest accuracy improvement (87%) among the previously failed samples. This result clearly demonstrates GPT-4o’s latent reasoning capabilities, significantly activated when explicitly guided through structured logical steps.
Similarly, the focus_logic prompt, explicitly mentioning the possibility of belief inconsistency and urging strict logical evaluation, substantially improved GPT-4o’s accuracy (76%). This finding suggests that GPT-4o is sensitive to explicit instructional framing and context-setting, effectively reducing its reliance on superficial plausibility heuristics when directed accordingly.
Conversely, informal prompts (casual) and simple instruction (basic) achieved extremely low recovery rates (3% and 5%, respectively). These prompts appear insufficient in addressing GPT-4o’s belief bias, indicating that the model defaults to common-sense heuristics without clear guidance.
The polite ({CJK}UTF8ipxm敬語, keigo) prompt yielded a moderate accuracy improvement (65%), suggesting linguistic politeness cues may moderately encourage GPT-4o to engage in deeper reflection or careful reasoning, though less effectively than explicitly logical or CoT framing.
Repeating this experiment with English-language prompts while keeping the Japanese data constant yielded a similar performance pattern, as illustrated in Figure 4. However, the accuracy gaps were less pronounced compared to the Japanese prompts. This narrower gap in performance likely stems from GPT-4o being extensively trained in English, enhancing its robustness to varied prompt styles. Nevertheless, this confirms the key insight: prompt design significantly influences GPT-4o’s performance on challenging tasks like belief-inconsistent syllogisms.
Figure 4: Prompt Type Accuracy on Error Samples (%) – Retest with English prompts
Table 2 provides detailed descriptions of each prompt type used in the error re-evaluation.
Table 2: Descriptions of prompt types used for re-evaluating GPT-4o errors.
Implications for Model Deployment
These findings demonstrate the substantial impact of prompt engineering on GPT-4o’s logical reasoning capabilities, particularly in overcoming belief bias. Although GPT-4o already performs strongly relative to other models, explicit prompting—such as instructing it to follow a structured reasoning process or clearly signaling the presence of belief-inconsistent content—markedly enhances its logical consistency. Consequently, when deploying LLMs, especially in contexts demanding precise and unbiased logical inference, strategic prompt design is crucial. Clear, structured instructions, emphasizing logical rigor and explicitly guiding step-by-step reasoning, can significantly mitigate intuitive biases inherent to the model, thus ensuring more reliable and accurate outcomes.
4.5 Logit Lens Analysis for Open Models
Based on the results in Table 1, we observe that the open model llm-jp-3-13b achieves competitive performance compared to commercial alternatives. To gain deeper insights into the model’s internal state during the belief-inconsistent syllogistic reasoning task, we conduct an analysis using the logit lens technique.
We perform this analysis on two models: llm-jp/llm-jp-3-13b and stockmark/stockmark-13b. Both are open, Japanese-optimized models of equal size (13B parameters), making them ideal for controlled comparison. Figure 5 presents a representative example of the logit lens output.
The key observation is the model’s behavior at the Conclusion_End token position. For stockmark/stockmark-13b, its prediction probability becomes high early in the network, suggesting a form of premature confidence. In contrast, llm-jp/llm-jp-3-13b shows a gradual increase in this signal across layers, indicating a delayed and layered buildup of reasoning. Additionally, the Premise1_End signal is notably weaker in stockmark, implying possible uncertainty in boundary detection between the premises and the conclusion.
(a) stockmark/stockmark-13b
(b) llm-jp/llm-jp-3-13b
Figure 5: Logit lens visualization on the belief-inconsistent syllogistic reasoning task.
5 Discussion
The results of our evaluation highlight important distinctions in the reasoning abilities of contemporary language models when confronted with belief-inconsistent syllogisms. GPT-4o’s leading performance—nearly 80% accuracy—suggests that advanced training strategies, possibly including large-scale instruction tuning and reasoning-augmented supervision, have enabled it to prioritize formal logic over commonsense intuition. Its effectiveness on a Japanese-language task, despite not being explicitly tailored for Japanese, underscores the strength of cross-lingual transfer when reasoning is properly grounded.
In contrast, the poor performance of models like Claude-3-opus raises concerns about reasoning suppression in alignment-optimized architectures. While such models may be calibrated to avoid generating implausible outputs, this safety-oriented behavior appears to interfere with their ability to recognize and validate logically correct but counterintuitive conclusions. Similarly, the moderate but consistent performance of Japanese-specialized models suggests that strong language modeling does not necessarily imply strong reasoning. These models handle Japanese fluently yet often fail to maintain deductive consistency, particularly when logic conflicts with expectation.
Another notable finding is the absence of a positive correlation between model size and reasoning performance. Larger models in our evaluation, such as llm-jp-3-172b, underperformed their smaller counterparts, indicating that scale alone is not sufficient. Rather, performance on belief-inconsistent reasoning tasks appears to depend more on the nature of pretraining data, task-aligned fine-tuning, and perhaps architectural inductive biases that support structured inference.
Overall, our results reinforce the view that logical reasoning remains a distinct capability in language models—one that must be explicitly tested and optimized for, particularly in high-stakes domains where trust in correctness cannot rest solely on fluency or apparent plausibility.
6 Limitations
While BIS Reasoning 1.0 reveals important insights into the reasoning capabilities of LLMs, this study has several limitations that should be considered. First, the evaluation focuses exclusively on syllogistic reasoning. While syllogisms offer a controlled and interpretable format, they represent only one class of logical reasoning. The results may not generalize to other reasoning forms such as causal inference, probabilistic reasoning, or multi-hop deductive chains.
Second, our evaluation relies on a single prompt design in a zero-shot setting. While this approach offers a consistent testbed, it may not fully capture the capabilities of models that perform better under more advanced prompting strategies or tailored task formulations. Prompt sensitivity remains an open variable, and performance may vary under alternative instructions or reasoning scaffolds.
Third, our scoring metric is binary and does not account for partially correct reasoning or near-misses. Models that correctly identify the logical structure but misword the conclusion, or those that reason correctly but fail to flag belief conflict, are treated the same as entirely incorrect responses.
In terms of model coverage, while we include both general-purpose and Japanese-specialized LLMs, our selection remains limited. Notably, open-source models outside of the LLM-JP and Stockmark ecosystems, as well as mid-sized multilingual models, are not represented in this evaluation.
Lastly, model capabilities are evolving rapidly. The results presented reflect the state of model behavior at a specific point in time (early 2025), and future model updates could yield different performance profiles.
7 Conclusion
We introduced BIS Reasoning 1.0, the first large-scale Japanese-language dataset explicitly focused on belief-inconsistent syllogistic reasoning. Through a comprehensive evaluation of state-of-the-art language models, we demonstrated that reasoning under belief conflict remains a fundamental weakness—even for models with native-level Japanese capabilities. Among the models tested, GPT-4o achieved the highest accuracy, showing that architectural and training sophistication can help mitigate (though not eliminate) this form of reasoning bias.
Our results underscore a critical insight: language fluency does not guarantee logical soundness. The performance gap between model families challenges the assumption that scale or localization alone ensures robustness, and reinforces the importance of domain-targeted benchmarks for high-stakes applications. BIS Reasoning 1.0 fills this gap by offering a linguistically natural, cognitively demanding benchmark designed to expose belief-driven reasoning failures.
This work lays the groundwork for future research on belief bias mitigation, logic-aware model design, and dataset extensions across new reasoning paradigms and specialized domains. As LLMs become more deeply embedded in fields such as law, healthcare, and policy, BIS Reasoning 1.0 provides a practical tool for stress-testing their logical fidelity in scenarios where correctness must outweigh intuition.
Acknowledgements
We would like to thank Yusuke Oda, Kiyomaru Hirokazu, Maki Matsuda, and Kouta Nakayama, among others, for their valuable input—ranging from ideas for evaluating large language models, suggestions that improved the literature review, to support in drafting annotation guidelines.